OpenStudy (anonymous):

find the sum of 1/5+ 3/8 =

OpenStudy (anonymous):

.575...if it helps?

OpenStudy (anonymous):

.4

OpenStudy (anonymous):

how'd u get tht

OpenStudy (anonymous):

on the calculator on my computer, i typed in 1/5+3/8 and it equaled .4

OpenStudy (anonymous):

ok but you needed to put (1/5)+(3/8) or else it will divide 1/5 and 3/8 then add them together. so i got .575

OpenStudy (anonymous):

are you still looking for the answer

OpenStudy (anonymous):

the answer is 21/40

OpenStudy (ybarrap):

1/5 + 3/8 find a common denominator 8/40 + 3*5/40 = (8 + 15)/40 = 23/40 = .575 Ans. 23/40

OpenStudy (anonymous):

wow i forgot to put it back in tht form good job ybarrap